OpenStudy (anonymous):

HELP ME PLEASE!!!!!!!!!!!!!!!!!! I will give u a medal Describe a situation that can be expressed with a two-step equation. Write the equation, solve it, and check the solution, showing all your work in word form

OpenStudy (godlovesme):

Hang on I'll help u :)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

HELP PLZ

OpenStudy (anonymous):

@sleepyjess

OpenStudy (godlovesme):

Srry for taking so long I was helping someone with long problem =0 Here: Joey's bikes rents bikes for $15 plus $5 per hour. Janice paid $30 to rent a bike. How many hours did she rent the bike Our equation is 5x+15=30

OpenStudy (anonymous):

30/15=2 ?

OpenStudy (godlovesme):

Ummm.. Let's c 5x+15-15 = 30-15 --> subtract 15 from both side 5x/5=15/5 --> divide both side by 5 X=3 --> cookies lol c:

OpenStudy (godlovesme):

To check plug 3 for x in the original equation and prove the answer is correct and u r done!! Hope that helped :D @divergentmockingjay19